OpenStudy (anonymous):

I don't understand about the vertical asymptote. Here is the problem: y = 1/( x – 4)– 2. How do I find a vertical asymptote?.

OpenStudy (anonymous):

a vertical asymtote occurs when y approches infinety. you will get y= inft. when: \[1/( x – 4)– 2\] = \[1/0\] so solve the equation for x where \[( x – 4)– 2=0\], will give you the x value at which you have a vertical asymptote. in this case at x= 6 you have a vertical asymptote. Do you get it? :)

OpenStudy (anonymous):

Simply put, let denominator = 0 => x = .. is vertical asymptote!

OpenStudy (anonymous):

I guess I don't understand. I keep looking, but I am just blank.

OpenStudy (anonymous):

I get that the denominator will be 0, so in this case I guess the x is the vertical asymptote. why is it not 4 instead of 6?

OpenStudy (anonymous):

because if x=4 you will not get the denominator to be zero. try and plug in x=4.. will that make y=1/0 ?

OpenStudy (anonymous):

could you work this out in steps so I can actually see it on paper? I'm sorry I am so dense.

OpenStudy (anonymous):

it is totally fine, check this out: http://www.youtube.com/watch?v=_qEOZNPce60

OpenStudy (anonymous):

Ok, Thanks for the insight. The video really helped me see what I needed to do. You are great.

OpenStudy (anonymous):

no problem, remember to give me a medal, just click best answer button :D